### v2

# Prosoluble subgroups of the profinite completion of the fundamental group of compact 33 33-manifolds

Abstract
We give a description of finitely generated prosoluble subgroups of the profinite completion of 33 33-manifold groups and virtually compact special groups.
MSC Classification: 20E08, 20E18.
Keywords: manifolds, fundamental groups, profinite groups, acylindrical actions.
1 Introduction
In recent years there has been a great deal of interest in detecting properties of the fundamental group $\pi_{1}(M)$ of a 33 33-manifold via its finite quotients, or more conceptually by its profinite completion. This motivates the study of the profinite completion $\widehat{\pi_{1}(M)}$ of the fundamental group of a 33 33-manifold. The study of $\widehat{\pi_{1}(M)}$ is based on the work of Henry Wilton and the second author (see [WZ17]) where it is proved that the standard splittings of 33 33-manifold groups as free product with amalgamation, HNN-extension and generally as a graph of groups (we refer to them as free constructions later) survive after profinite completion of $\pi_{1}(M)$. Namely $\widehat{\pi_{1}(M)}$ admits the (profinite) Kneser–Milnor decomposition and then the (profinite) JSJ-decomposition, and for hyperbolic manifolds with cusps (after passing to a finite sheeted cover) there is a suitable hierarchy in which the corresponding actions on profinite trees are profinitely acylindrical.
At this point one can use the profinite analogue of Bass-Serre theory for groups acting on trees. Note, however, that this theory does not have the full strength of its classical original. The main theorem of Bass-Serre theory (that describes the subgroups of free constructions) asserts that a group $G$ acting on a tree $T$ is the fundamental group of a graph of groups $(\mathcal{G},G\backslash T,D)$, where $D$ is a maximal subtree of $G\backslash T$ and $\mathcal{G}$ consists of edge and vertex stabilizers of a connected transversal of $G\backslash T$. This does not hold in the profinite case. As a consequence, subgroup theorems do not hold, in general, for profinite groups and even for free profinite products, i.e., the profinite version of the Kurosh subgroup theorem is not valid. This implies that it is reasonable to study subgroup structure of free constructions for important subclasses of profinite groups. The most important subclasses of profinite groups are prosoluble and pro-$p$ as they play the same role as finite $p$-groups and soluble groups in finite group theory. In particular, this applies to the profinite completion of 33 33-manifold groups.
Such a study for pro-$p$ subgroups was performed in [WZ18], where the structure of finitely generated pro-$p$ subgroups of profinite completions of the fundamental groups of a 33 33-manifold was described. The objective of this paper is to study prosoluble subgroups of $\widehat{\pi_{1}(M)}$.
We begin with the fundamental group of a compact hyperbolic 33 33-manifold.
Theorem A.
Let $\pi_{1}(M)$ be the fundamental group of a compact hyperbolic 33 33-manifold $M$ and $H$ a finitely generated prosoluble subgroup of the profinite completion of $\pi_{1}(M)$. Then $H$ is projective, i.e., of cohomological dimension 11 11.
Note that the Schreier theorem does not hold for profinite groups, i.e., subgroups of free profinite groups, in general, are not free but projective. So Theorem A tells that prosoluble subgroups of $\widehat{\pi_{1}(M)}$ are isomorphic to subgroups of a free profinite group.
The proof of Theorem A uses the dramatic developments of Agol ([Ago13]), Kahn-Markovic ([KM12]) and Wise ([Wis21]), whose works implies that the fundamental groups of closed hyperbolic 33 33-manifolds are also fundamental groups of compact, virtually special cube complexes (see [AFW15] for a summary of these developments); such groups are called virtually compact special. Virtually special groups are originated in the works of Wise. These groups have a central role in modern geometric group theory and Haglund and Wise showed that many hyperbolic groups are virtually special such as hyperbolic Coxeter groups and one relator groups with torsion.
The next theorem describes torsion-free finitely generated prosoluble subgroup of virtually special groups.
Theorem B.
Let $G$ be a torsion-free, hyperbolic, virtually special group. Any finitely generated prosoluble subgroup $H$ of the profinite completion of $G$ is projective.
Using that the standard cocompact arithmetic subgroups of $\operatorname{SO}(n,1)$ are virtually compact special ([BHW11, Theorem 1.10]) we deduce the following
Corollary.
Prosoluble subgroups of the profinite completion of standard cocompact arithmetic subgroups of $\operatorname{SO}(n,1)$ are projective.
The next objective of this paper is to extend these results to relatively hyperbolic virtually compact special groups, in particular, to the fundamental hyperbolic 33 33-manifolds with cusps. It is based on relatively hyperbolic analogs of Wise’s Quasiconvex Hierarchy Theorem proved by Eduardo Einstein [Ein19] and the fact that it is preserved by the profinite completion established in [Zal23a].
Theorem C.
Let $G$ be a torsion-free virtually compact special toral relatively hyperbolic group and $H$ a finitely generated prosoluble subgroup of the profinite completion $\widehat{G}$ of $G$. Then one of the following statements holds:
  1. (i)
    $H$ is virtually a subgroup of a free prosoluble product of abelian pro-$p$ groups;
  2. (ii)
    $H$ is a virtually abelian group.
Using the [WZ18, Theorem 4.3] for 33 33-manifolds with cusps we deduce the following:
Theorem D.
Let $M$ be a hyperbolic 33 33-manifolds with cusps and $H$ a finitely generated prosoluble subgroup of the profinite completion of $\pi_{1}(M)$. Then one of the following statements holds:
  1. (i)
    $H$ is virtually a subgroup of a free prosoluble product of abelian pro-$p$ groups;
  2. (ii)
    $H$ is a virtually abelian.
Finally, analyzing then $\widehat{\pi_{1}(M)}$ for Thurston’s geometries case-by-case, we obtain the classification of finitely generated prosoluble subgroups of $\widehat{\pi_{1}(M)}$ for an arbitrary compact, orientable 33 33-manifold.
Theorem E.
Let $M$ be a compact orientable 33 33-manifold. If $H$ is a finitely generated prosoluble subgroup of $\widehat{\pi_{1}(M)}$, then one of the following statements holds:
  1. (i)
    $H$ is isomorphic to a subgroup of a free prosoluble product of the pro-$p$ groups from the following list of isomorphism types:
    1. (1)
      For $p>3$: $C_{p}$; $\mathbb{Z}_{p}$; $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$; the pro-$p$ completion of $(\mathbb{Z}\times\mathbb{Z})\rtimes\mathbb{Z}$ and the pro-$p$ completion of a residually-$p$ fundamental group of a non-compact Seifert fibred manifold with hyperbolic base of orbifold;
    2. (2)
      For $p=3$: in addition to the list of (1) we have a torsion-free extension of $\mathbb{Z}_{3}\times\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ by $C_{3}$;
    3. (3)
      For $p=2$: in addition to the list of (1) we have $C_{2^{m}}$; $D_{2^{k}}$; $Q_{2^{n}}$; $\mathbb{Z}_{2}\rtimes C_{2}$; the torsion-free extensions of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ by one of $C_{2}$, $C_{4}$, $C_{8}$, $D_{2}$, $D_{4}$, $D_{8}$, $Q_{16}$; the pro-22 22 extension of the Klein-bottle group $\mathbb{Z}\rtimes\mathbb{Z}$; the pro-22 22 completion of all torsion-free extensions of a soluble group $(\mathbb{Z}\rtimes\mathbb{Z})\rtimes\mathbb{Z}$ with a group of order at most 22 22;
  2. (ii)
    $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group where $C$ is a finite cyclic group and $\pi$ is set of primes;
  3. (iii)
    $H$ is a subgroup of a central extension of $D_{2n}$, a tetrahedral group $T$, a octahedral group $O$ or a icosahedral group $I$ by a cyclic group of even order;
  4. (iv)
    $H$ is a subgroup of the profinite completion of a 33 33-dimensional Bieberbach group $B$ (i.e, $B$ is torsion-free virtually $\mathbb{Z}^{3}$);
  5. (v)
    $H$ is a subgroup of the profinite completion of a group extension of $\mathbb{Z}^{2}\rtimes_{A}\mathbb{Z}$ by $C$ where $C$ is trivial or $C_{2}$ and $A$ is an Anosov matrix;
  6. (vi)
    $H$ is an extension of a torsion-free procyclic group ^$\widehat{\mathbb{Z}}_{\sigma}$ by a subgroup $H_{0}$ of a finite free prosoluble product of finite cyclic $p$-groups with with $H_{0}$ acting either trivially on ^$\widehat{\mathbb{Z}}_{\sigma}$ or by inversion.
In this theorem we are denoting by $T$ the tetrahedral group of 1212 1212 rotational symmetries of a tetrahedron, by $O$ the octahedral group of 2424 2424 rotational symmetries of a cube or of an octahedron, by $I$ the icosahedral group of 6060 6060 rotational symmetries of a dodecahedron or an icosahedron. An Anosov matrix is a matrix $A\in\operatorname{GL}_{2}(\mathbb{Z})$ determined by an Anosov homeomorphism of the torus satisfying the following conditions $\operatorname{tr}(A)>2$ if $\det(A)=1$ and $\operatorname{tr}(A)\neq 0$ if $\det(A)=-1$ (see [GM23, Section 2]). Also, we use the notation ^$\widehat{\mathbb{Z}}_{\alpha}=\prod_{p\in\alpha}\mathbb{Z}_{p}$ where $\alpha$ is a set of primes.
To prove Theorem E we make heavy use of the profinite version of Bass-Serre theory. In particular, we prove the following general theorem of independent interest:.
Theorem F.
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups, $G$ its fundamental group. If $H$ is a finitely generated prosoluble subgroup of $G$, then $H$ embeds into a free profinite product $\coprod_{v\in\Gamma}H_{v}$, where $H_{v}=H\cap\mathcal{G}(v)^{g}$ and $g\in G$. Moreover, if $H\neq H_{v}$ for some $v$, then either $H_{v}$ are finitely generated pro-$p$ or $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group (i.e., $C$ is cyclic, $|C|$ divides $p-1$ for any $p\in\pi$ and $[c,z]\neq 1$ for all $1\neq c\in C,0\neq z\in\widehat{\mathbb{Z}}_{\pi}$).
The celebrated Sela theorem (see [Sel97, Theorem 3.2]) states that a finitely generated group is $k$-acylindrically accessible, i.e., if it acts on a tree $T$ then $G\backslash T$ is finite and, in fact, bounded in terms of $k$ and the minimal number of generators $d(G)$. This is equivalent to the statement that $G$ has finitely many, bounded, non-trivial vertices stabilizers up to conjugation. These statements are not equivalent however for profinite groups (one can see it using Proposition 2.15). Thus the correct profinite analog of accessibility is a bound for the number of non-trivial vertex stabilizers up to conjugation. Using Theorem F we prove an analogue of Sela’s theorem for prosoluble groups.
Theorem G.
Let $G$ be a finitely generated prosoluble group acting on simply connected profinite graph $\Gamma$111simply connected means $\pi_{1}(\Gamma)=1$ and this is not the same as acyclicity in the profinite case (cf. Definition 2.4).. Then the number of vertex stabilizers of $G$, up to conjugation, is at most $d(G)$.
This paper is organized as follows. In Section 2 we introduce the basic ideas and notations of profinite Bass-Serre theory as well as prove auxiliary results. In Section 3, we study pro-$\mathcal{C}$ groups acting $k$-acylindrically on a profinite simply connected graph. Closed pro-$\mathcal{C}$ subgroups of the profinite fundamental groups of a finite graph of profinite groups with finite edge groups are studied in Section 4 and Section 5 dedicated to finitely generated prosoluble subgroups of the fundamental group of $k$-acylindrical finite graph of profinite groups. Section 6 is dedicated to prosoluble subgroups of virtually special groups. We use it in Section 7 to prove Theorem D. In Section 8 we deal with Thurston’s geometries of a compact 33 33-manifold in order to prove Theorem E.
When we deal with profinite groups our subgroups are assumed to be closed, homomorphisms to be continuous, finitely generated means topologically finitely generated. Throughout the paper we denote by $\mathcal{C}$ a variety closed to extensions according to the terminology of [RZ10]. We shall restate the results of the introduction inside of the section for easier reading.
2 Finite graphs of pro-$\mathcal{C}$ groups and its fundamental group
In this section we will introduce the basic definitions and notations of profinite Bass-Serre theory used in the paper. This section is entirely based on [Rib17].
2.1 Finite graphs of pro-$\mathcal{C}$ groups
Definition 2.1 (Profinite graphs).
A profinite graph is a profinite space $\Gamma$ with a distinguished non-empty subset $V(\Gamma)$ and two continuous maps $d_{i}:\Gamma\to V(\Gamma)$, $i=0,1$, which are identity on $V$.
The subset $V(\Gamma)$ is the vertex-set of $\Gamma$ and $E(\Gamma)=\Gamma-V$ is the edge-set of $\Gamma$. If $e\in E(\Gamma)$, then we have
$d_{0}(e)$$e$$d_{1}(e)$
We call $d_{0}(e)$ and $d_{1}(e)$, respectively, by initial and terminal vertex. The maps $d_{0},d_{1}$ are the incidence maps.
Definition 2.2.
A morphism $\varphi:\Gamma\to\Delta$ of profinite graphs is a continuous map with $\varphi d_{i}=d_{i}\varphi$ for $i=0,1$.
Definition 2.3.
A profinite graph $\Gamma$ is connected if given any morphism of profinite graphs $\varphi$ from $\Gamma$ onto a finite graph $\Sigma$, then $\Sigma$ is connected as an abstract graph.
Definition 2.4 (Simply connected profinite graphs).
Let $G$ be a pro-$\mathcal{C}$ group acting freely on a profinite graph $\Gamma$. The natural epimorphism $\zeta:\Gamma\to G\backslash\Gamma$ is called a Galois $\mathcal{C}$-covering of $G\backslash\Gamma$ and the associated group is denoted by $G=G(\zeta)$. If $\Gamma$ does not have non-trivial Galois $\mathcal{C}$-coverings, then $\zeta$ is called universal $\mathcal{C}$-covering and, in this case, $\pi_{1}(\Gamma)=G(\Gamma|\Delta)$ is the pro-$\mathcal{C}$ fundamental group of $\Gamma$. If $\pi_{1}(\Gamma)=1$ then we say that $\Gamma$ is $\mathcal{C}$-simply connected.
Let $\Gamma$ be a profinite graph. Define $E^{\ast}(\Gamma)=\Gamma\backslash V(\Gamma)$ and ^$\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}$ as the pro-$\mathcal{C}$ completion of $\mathbb{Z}$.
Definition 2.5 (Profinite trees).
We say that $\Gamma$ is a pro-$\mathcal{C}$ tree if the sequence
$0\displaystyle 0\buildrel\over{\longrightarrow}[[\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}(E^{\ast}(\Gamma),\ast)]]\buildrel D\over{\longrightarrow}[[\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}(V(\Gamma))]]\buildrel\varepsilon\over{\longrightarrow}\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}\buildrel\over{\longrightarrow}$ 0
is exact, where $d(\bar{e})=d_{1}(e)-d_{0}(e)$ for the image $\bar{e}$ of $e\in E(\Gamma)$ in $E^{\ast}(\Gamma)$ with $d(\ast)=0$ and $\varepsilon(v)=1$ for $v\in V(\Gamma)$.
In [Rib17, Section 3.10] we can see that `````` `every simply connected profinite graph is a profinite tree""", `````` `if $\mathcal{C}$ is the variety of all finite soluble groups, $\mathcal{C}$-simply connected and $\mathcal{C}$-tree are equivalent""".
Also, if we define
$T^{g}=\{m\in T:gm=m\}$
to be the set of $g$-fixed points for a pro-$\mathcal{C}$ tree $T$, then $T^{g}$ is again a $\mathcal{C}$-tree (see [Rib17, Theorem 4.1.5]).
Definition 2.6.
Let $\Gamma$ be a pro-$\mathcal{C}$ tree. Given the vertices $v,w\in\Gamma$ we can consider $[v,w]$ as the intersection of all subtrees of $\Gamma$ containing $v$ and $w$. We call $[v,w]$ of geodesic. If $[v,w]$ is a reduced path $p_{v,w}$ from $v$ to $w$, the length $\ell([v,w])$ of $[v,w]$ is the number of edges in $p_{v,w}$.
Definition 2.7 (Finite graphs of profinite groups).
A finite graph of pro-$\mathcal{C}$ groups is a pair $(\mathcal{G},\Gamma)$ where $\Gamma$ is a connected finite graph, $\mathcal{G}=\dot{\bigcup_{m\in\Gamma}}\mathcal{G}(m)$ with $\mathcal{G}(m)$ a pro-$\mathcal{C}$ group for each $m$ and there are continuous monomorphisms $\partial_{i}:\mathcal{G}(e)\to\mathcal{G}(d_{i}(e))$ for each edge $e\in E(\Gamma)$.
Definition 2.8.
A morphism
$\underline{\alpha}=(\alpha,\alpha^{\prime}):(\mathcal{G},T)\to(\mathcal{G}^{\prime},T^{\prime})$
of graphs of pro-$\mathcal{C}$ groups consists of a pair of continuous maps $\alpha:\mathcal{G}\to\mathcal{G}^{\prime}$ and $\alpha^{\prime}:T\to T^{\prime}$ such that the diagram
[…]
is commutative. We say that $\underline{\alpha}$ is a monomorphism if $\alpha$ and $\alpha^{\prime}$ are injective.
Definition 2.9.
A finite graph of pro-$\mathcal{C}$ groups $(\mathcal{G},\Gamma)$ is reduced if for every edge $e$ which is not a loop (that is, $d_{0}(e)\neq d_{1}(e)$), neither $\partial_{0}:\mathcal{G}(e)\to\mathcal{G}(d_{0}(e))$ nor $\partial_{1}:\mathcal{G}(e)\to\mathcal{G}(d_{1}(e))$ is an isomorphism. We say that an edge $e$ is fictitious if $e$ is not a loop and $\partial_{0}$ or $\partial_{1}$ is an isomorphism.
Given a finite graph of groups with a fictitious edges $e_{1},...,e_{n}$ we can collapse this edges transforming it in a reduced finite graph of groups with the same fundamental group (see [CZ23, Remark 2.8]).
2.2 The fundamental group of a finite graph of pro-$\mathcal{C}$ groups
For a totally new construction of the fundamental group of a profinite graph of pro-$\mathcal{C}$ groups using an approach similar to the abstract case, the reader should check [AZ22].
Definition 2.10 (The fundamental group of a finite graph of profinite groups).
Let $(\mathcal{G},\Gamma)$ be a finite graph of pro-$\mathcal{C}$ groups and $T$ be a maximal subtree of $\Gamma$. The pro-$\mathcal{C}$ fundamental group of $(\mathcal{G},\Gamma)$ is a group $G$ with continuous homomorphisms
$\nu_{m}:\mathcal{G}(m)\to G$
and continuous maps $E(\Gamma)\to G$ denoted by $e\mapsto t_{e}$ such that $t_{e}=1$ if $e\in E(T)$ and
$(\nu_{d_{0}(e)}\partial_{0})(x)=t_{e}(\nu_{d_{1}(e)}\partial_{1})(x)t_{e}^{-1}$
for all $x\in\mathcal{G}(e)$ and $e\in E(\Gamma)$ satisfying the following universal property: if $H$ is a pro-$\mathcal{C}$ group, $\beta_{m}:\mathcal{G}(m)\to H$ are continuous homomorphisms, $e\to s_{e}$ ($e\in E(\Gamma)$) is a map with $s_{e}=1$ if $e\in E(T)$ and
$(\beta_{d_{0}(e)}\partial_{0})(x)=t_{e}(\beta_{d_{1}(e)}\partial_{1})(x)t_{e}^{-1}$
for all $x\in\mathcal{G}(e)$ and $e\in E(\Gamma)$, then there is a unique continuous homomorphism $\delta:G\to H$ such that $\delta(t_{e})=s_{e}$ and $\delta\nu_{m}=\beta_{m}$ for each $m\in\Gamma$.
We will denote the fundamental group of $(\mathcal{G},\Gamma)$ by
$G=\Pi_{1}(\mathcal{G},\Gamma).$
The reader can find proofs of existence and uniqueness of fundamental groups in [Rib17].
The basic free constructions such as free profinite groups, free profinite products, free profinite products with amalgamation and $\operatorname{HNN}$-extensions are the most basic examples of fundamental groups of finite graphs of groups (see [Rib17, Section 6.2]).
The following lemma will be useful:
Lemma 2.11.
Let $(\mathcal{G},\Gamma)$ be an injective profinite graph of pro-$\mathcal{C}$ groups and $\Delta$ a connected subgraph of $\Gamma$. Then $\Pi_{1}(\mathcal{G},\Delta)$ is embedded in $\Pi_{1}(\mathcal{G},\Gamma)$.
Proof.
First assume that $(\mathcal{G},\Gamma)$ is a finite graph of finite groups. Note that $\Pi_{1}(\mathcal{G},\Gamma)$ and $\Pi_{1}(\mathcal{G},\Delta)$ are pro-$\mathcal{C}$ completions of the abstract fundamental groups ${\Pi_{1}}^{\text{abs}}(\mathcal{G},\Gamma)$ and ${\Pi_{1}}^{\text{abs}}(\mathcal{G},\Delta)$, respectively. Consider the collapse $\Gamma\to\Gamma/\Delta$ and the graph of groups $(\mathcal{G}_{\Delta},\Gamma/\Delta)$ defining $\mathcal{G}_{\Delta}(m)={\Pi_{1}}^{\text{abs}}(\mathcal{G},\Delta)$ if $m$ is the image of $\Delta$ and $\mathcal{G}_{\Delta}(m)=\mathcal{G}(m)$ for the others $m$. Since the edge groups of $(\mathcal{G}_{\Delta},\Gamma/\Delta)$ are finite, the pro-$\mathcal{C}$ topology of ${\Pi_{1}}^{\text{abs}}(\mathcal{G}_{\Delta},\Gamma/\Delta)={\Pi_{1}}^{\text{abs}}(\mathcal{G},\Gamma)$ induces the full pro-$\mathcal{C}$ topology on ${\Pi_{1}}^{\text{abs}}(\mathcal{G},\Delta)$ (see [Rib17, Proposition 6.5.5]). Then by [Rib17, Proposition 6.5.3] the graph of groups $(\overline{\mathcal{G}}_{\Delta},\Gamma/\Delta)$ of the pro-$\mathcal{C}$ completions of vertex and edge groups is injective. It follows that the pro-$\mathcal{C}$ completion of ${\Pi_{1}}^{\text{abs}}(\mathcal{G},\Delta)$ embeds in the pro-$\mathcal{C}$ completion of ${\Pi_{1}}^{\text{abs}}(\mathcal{G},\Gamma)$, as desired.
For the general case, given any open normal subgroup $U$ of $G$, we have (see [AZ22, Proposition 2.15 and Theorem 3.9])
$G_{U}=G/\tilde{U}=\Pi_{1}(\mathcal{G}_{U},\Gamma)$
such that
[…]
and
[…]
Since the diagram
$\Pi_{1}(\mathcal{G},\Gamma)$$\Pi_{1}(\mathcal{G},\Delta)$$\Pi_{1}(\mathcal{G}_{U},\Gamma)$$\Pi_{1}(\mathcal{G}_{U},\Delta)$
commutes, we get $\Pi_{1}(\mathcal{G},\Delta)\leq\Pi_{1}(\mathcal{G},\Gamma)$. ∎
With standard arguments and can deduce the following:
Proposition 2.12.
Let $G$ be the fundamental group of a finite reduced graph of profinite groups with at least one edge. If $G$ is not virtually cyclic, then $G$ contains a non-abelian free profinite group. In particular, if $\mathcal{C}$ does not contains all finite groups, then $G$ is not pro-$\mathcal{C}$.
Proof.
Let $U$ be an open normal subgroup and
$\tilde{U}=\left\langle U\cap\mathcal{G}(v)^{g}:v\in V(\Gamma),g\in G\right\rangle.$
[…]
Definition 2.13 (Standard graph).
Given a finite graph of pro-$\mathcal{C}$ groups $(\mathcal{G},\Gamma)$ with fundamental group $G$, the correspondent standard graph is
$S=S(G)=\dot{\bigcup_{m\in\Gamma}}G/\mathcal{G}(m).$
The vertices of $S$ are the cosets $g\mathcal{G}(v)$ with $v\in V(\Gamma)$ and $g\in G$; the edges of $S$ are the cosets $g\mathcal{G}(e)$ with $e\in E(\Gamma)$ and $g\in G$; the incidence maps of $S$ are given by
$d_{0}(g\mathcal{G}(e))=g\mathcal{G}(d_{0}(e)),\quad d_{1}(g\mathcal{G}(e))=gt_{e}\mathcal{G}(d_{1}(e))$
with $e\in E(\Gamma)$ and $t_{e}=1$ if $e\in E(T)$.
The fact that $S$ is indeed a profinite graph as well as the basic properties of $S$ are explained in [Rib17, Section 6.3]. In particular, it is proved in [Rib17, Theorem 6.3.5] that $S$ is $\mathcal{C}$-simply connected and so is a pro-$\mathcal{C}$ tree. Besides, $G$ acts continuously on $S$ with $G\backslash S\simeq\Gamma$.
2.3 Auxiliary Results
In this subsection we will prove some facts relevant to the general theory of profinite groups acting on profinite graphs. The results here will be helpful to show that $k$-acylindrical actions are well-behaved under taking quotients in the next section.
Lemma 2.14.
Let $G$ be a profinite group acting on a simply connected profinite graph $\Gamma$ and $\Sigma$ a finite connected subgraph of $\Gamma$. Then $H=\left\langle G_{v}\mid v\in V(\Sigma)\right\rangle$ is the fundamental group of a finite tree of pro-$\mathcal{C}$ groups $\Pi_{1}(\mathcal{H},\Delta)$ whose edge and vertex groups are stabilizers of the corresponding edges and vertices of a connected transversal of $\Delta=H\backslash H\Sigma$ in $T$.
Proof.
Put $T=H\Sigma$. Since the quotient of $T$ modulo every open subgroup $U$ of $H$ is connected, $T$ is connected. By [Rib17, Proposition 3.7.3], $\Sigma$ is simply connected. By [Rib17, Proposition 3.9.2], $\Delta=H\backslash T$ is simply connected and so is a finite tree. By [Rib17, Theorem 6.6.1], $H=\Pi_{1}(\mathcal{H},\Delta)$ whose edge and vertex groups are stabilizers of the corresponding edges and vertices of a connected transversal of $\Delta$ in $T$. ∎
Proposition 2.15.
Let $G$ be a profinite group acting on a simply connected profinite graph $\Gamma$ and let $e$ be an edge of $S$. Suppose the stabilizers of the vertices $v$, $w$ of the edge $e$ satisfy the following condition: $|G_{v}:G_{e}|+|G_{w}:G_{e}|>4$. Then the group $H=\left\langle G_{v},G_{w}\right\rangle$ is a free profinite amalgamated product $G_{v}\amalg_{G_{e}}G_{w}$. In particular, if $\mathcal{C}$ does not contains all finite groups, then $H$ is not pro-$\mathcal{C}$.
Proof.
By Lemma 2.14
$H=\Pi_{1}(\mathcal{H},H\backslash D)\simeq G_{v}\amalg_{G_{e}}G_{w}.$
Then, by Proposition 2.12, $H$ contains a non-abelian free profinite group and so cannot be pro-$\mathcal{C}$. ∎
The previous result does not holds if there is some $e$ such that $|G_{v}:G_{e}|=|G_{w}:G_{e}|=2$. When it occurs we will say that $\left\langle G_{v},G_{e},G_{w}\right\rangle$ is of dihedral type or that the edge $e$ generates a dihedral type group.
Corollary 2.16.
Let $\mathcal{C}$ be a variety does not containing all finite groups and $G$ a pro-$\mathcal{C}$ group acting on a simply connected profinite graph $\Gamma$. Suppose that $G$ has no edges $e$ generating a dihedral type group. Then for every edge $e$ with vertices $v,w$ and non-trivial edge stabilizer, we have $G_{v}=G_{e}$ or $G_{w}=G_{e}$.
Proof.
Suppose that, for a non-trivial edge stabilizer $e$ with vertices $v$ and $w$, we have $G_{v}$ and $G_{w}$ both different from $G_{e}$. Note that the index of $G_{e}$ in $G_{v}$ or in $G_{w}$ is greater than 22 22. By Proposition 2.15, $H=\left\langle G_{v},G_{w}\right\rangle$ is a not a pro-$\mathcal{C}$ subgroup of $G$, a contradiction. ∎
Lemma 2.17.
Let $G=G_{1}\amalg_{H}G_{2}$ be a splitting of a profinite group $G$ as an amalgamated free profinite product of profinite groups $G_{1},G_{2}$ and $H_{1}\leq G_{1},H_{2}\leq G_{2}$ be subgroups such that $H_{1}\cap H\leq U\geq H\cap H_{2}$ for some open normal subgroup $U$ of $G$. Then $\langle H_{1},H_{2}\rangle=L_{1}\amalg_{K}L_{2}$ with $L_{1}\leq H_{1}U,L_{2}\leq H_{2}U,K\leq HU$.
Proof.
By [ZM89, Proposition 4.4], $\langle H_{1},H_{2}\rangle=L_{1}\amalg_{K}L_{2}$ with $L_{1}\leq G_{1}$, $L_{2}\leq G_{2}$, $K\leq H$. Recall that the fundamental group $\Pi_{1}(\mathcal{G},\Gamma)$ of a finite graph of profinite groups is a completion of the abstract fundamental group ${\Pi_{1}}^{\text{abs}}(\mathcal{G},\Gamma)$ of the same graph of groups with respect to to the topology generated by the finite index normal subgroups $N$ of ${\Pi_{1}}^{\text{abs}}(\mathcal{G},\Gamma)$ such that $N\cap\mathcal{G}(v)$ is open in $\mathcal{G}(v)$. Note that $G$ is the fundamental group of a finite graph of profinite groups. We can consider $U\cap G_{1}$ and $U\cap G_{2}$ to apply [CZ23, Lemma 3.1] on ${G}^{\text{abs}}=G_{1}\ast_{H}G_{2}$ so we have ${\left\langle H_{1},H_{2}\right\rangle}^{\text{abs}}=L_{1}\ast_{K}L_{2}$ where $L_{1}\leq H_{1}(U\cap G_{1})$, $L_{2}\leq H_{2}(U\cap G_{2})$ and $K\leq H\cap U$. Then $\left\langle H_{1},H_{2}\right\rangle=\overline{L}_{1}\amalg_{K}\overline{L}_{2}$ with $\overline{L}_{1}\leq H_{1}U$, $\overline{L}_{2}\leq H_{2}U$ and $K\leq HU$. ∎
Corollary 2.18.
Let $G=G_{1}\amalg_{H}G_{2}$ be a splitting of the profinite group $G$ as an amalgamated free profinite product of profinite groups $G_{1},G_{2}$ and $H_{1}\leq G_{1},H_{2}\leq G_{2}$ be subgroups such that $H_{1}\cap H=1=H_{2}\cap H$. Then $\langle H_{1},H_{2}\rangle=H_{1}\amalg H_{2}$.
Proof.
Since $U$ in the preceding lemma is arbitrary, the result follows. ∎
2.4 Relatively projective groups
As was already mentioned in the introduction, the profinite analogue of the Kurosh Subgroup Theorem does not hold for free products of profinite groups and so the algebraic structure of closed subgroups of free profinite products is not clear. To compensate this, Dan Haran (see [Har87]) introduced the notion of relatively projective profinite groups and proved that a subgroup $H$ of a free profinite product $\coprod_{x\in X}G_{x}$ is projective relatively to the family $\{H\cap G_{x}^{g}\mid x\in X,g\in G\}$. We shall use a slightly more restricted notion here from [HZ23] as it better fits into the profinite version of Bass-Serre theory and suffices for our purpose.
Definition 2.19.
A group pile (or just pile) is a pair $\mathbf{G}=(G,T)$ consisting of a profinite group $G$, a profinite space $T$ and a continuous action of $G$ on $T$.
Definition 2.20.
A morphism of groups piles $\alpha:(G,T)\to(G^{\prime},T^{\prime})$ consists of a group homomorphism $\alpha:G\to G^{\prime}$ and a continuous map $\alpha:T\to T^{\prime}$ such that $\alpha(t^{g})=\alpha(t)^{\alpha(g)}$ for all $t\in T$ and $g\in G$. The above morphism is an epimorphism if $\alpha(G)=G^{\prime}$, $\alpha(T)=T^{\prime}$ and for every $t^{\prime}\in T^{\prime}$ there is $t\in T$ such that $\alpha(t)=t^{\prime}$ and $\alpha(G_{t})={G^{\prime}}_{t^{\prime}}$. It is rigid if $\alpha$ maps $G_{t}$ isomorphically onto ${G^{\prime}}_{\alpha(t)}$, for every $t\in T$, and the induced map of the orbit spaces $G\backslash T\to G^{\prime}\backslash T^{\prime}$ is a homeomorphism.
Definition 2.21.
An embedding problem for a pile $\mathbf{G}$ is a pair $(\varphi,\alpha)$ where $\varphi:\mathbf{G}\to\mathbf{H}$ and $\alpha:\mathbf{G}^{\prime}\to\mathbf{H}$ are morphism of group piles such that $\alpha$ is an epimorphism. It is finite if $\mathbf{G}^{\prime}$ is finite. It is rigid if $\alpha$ is rigid. A solution to the embedding problem is a morphism $\gamma:\mathbf{G}\to\mathbf{G}^{\prime}$ such that $\alpha\circ\gamma=\varphi$.
It means that the diagram
$(G^{\prime},T^{\prime})$$(H,X)$$(G,T)$$\gamma$$\alpha$$\varphi$
is commutative.
Definition 2.22.
A pile $\mathbf{G}$ is $\mathcal{C}$-projective if every finite rigid $\mathcal{C}$-embedding problem for $\mathbf{G}$ has a solution. A group $G$ is $\mathcal{G}$-projective if every finite $\mathcal{C}$-embedding problem has a weak solution.
It follows from [Zal23b, Remark 2.4] that if $\mathbf{G}$ is projective then $G$ is $\mathcal{G}$-projective.
We will make use of the following result:
Theorem 2.23.
[Zal23b, Theorem 4.1]
  1. (i)
    Let $G$ be a prosoluble group acting on a profinite tree $D$ with trivial edge stabilizers. Then $(G,V(D))$ is $\mathcal{S}$-projective, where $\mathcal{S}$ is the class of all finite soluble groups.
  2. (ii)
    Let $\mathcal{C}$ be a class of finite groups closed for subgroups, quotients and extensions. Let $(G,T)$ be a projective pro-$\mathcal{C}$ pile. Suppose that exists a continuous section $\sigma:G\backslash T\to T$. Then $G$ acts on a pro-$\mathcal{C}$ tree with trivial edge stabilizers and non-trivial vertex stabilizers being stabilizers of points in $T$.
We also need of the following proposition:
Proposition 2.24.
[Zal23b, Proposition 2.5] Let $\mathcal{C}$ be a class of finite groups closed for subgroups. quotients and extensions. Let $(G,T)$ be a $\mathcal{C}$-projective pile such that there exists a continuous section $\sigma:G\backslash T\to T$. Then $G$ embeds into a free pro-$\mathcal{C}$ product
$\coprod_{s\in Im(\sigma)}G_{s}\amalg F$
where $F$ is a free pro-$\mathcal{C}$ group of rank $d(G)$, the minimal number of generators of $G$. Moreover, every $G_{t}$ is of the form $G_{s}^{g}\cap G$ for some $g\in\coprod_{s\in Im(\sigma)}G_{s}\amalg F$, $s\in Im(\sigma).$
3 Acylindrical actions and prosoluble groups
In this section we prove a generalization of [WZ17, Lemma 11.2]. This result allows us to deduce that if $G$ acts $k$-acylindrically on a profinite simply connected graph $T$, then for any closed normal subgroups $N$ generated by vertex stabilizers the quotient group $G/N$ acts $k$-acylindrically on $N\backslash T$.
Definition 3.1.
An action of a profinite group $G$ on a profinite graph $T$ is called $k$-acylindrical if the stabilizer of any geodesic in $T$ of length greater than $k$ is trivial.
This definition is sometimes stated in the following way:
Definition 3.2.
An action of a profinite group $G$ on a profinite tree $T$ is called $k$-acylindrical if, for every nontrivial $g\in G$, the subtree of fixed points
$T^{g}=\{m\in T:gm=m\}.$
has diameter at most $k$.
It is easy to see that both definitions are equivalent: suppose that Definition 3.1 holds and there is a path $[v,w]$ with length greater than $k$ in $T^{g}$ for some nontrivial $g\in G$. Then $g\in G_{[v,w]}$, a contradiction. Conversely, if $T^{g}$ has diameter at most $k$ for every nontrivial $g\in G$ and $[v,w]$ is a geodesic of length greater than $k$. If $g\in G_{[v,w]}$, then $gv=v$ and $gw=w$ so that $[v,w]\in T^{g}$, a contradiction.
Definition 3.3.
We say that a finite graph of profinite groups $(\mathcal{G},\Gamma)$ is $k$-acylindrical if its fundamental group acts $k$-acylindrically on its standard pro-$\mathcal{C}$ tree.
Proposition 3.4.
Let $(\mathcal{G},\Gamma)$ be a $k$-acylindrical finite graph of profinite groups and $G$ its fundamental group. Then there is no edge $e$ such that
$|G_{d_{0}(e)}:G_{e}|=2=|G_{d_{1}(e)}:G_{e}|$
with $G_{e}\neq 1$.
Proof.
Assume that there is at least one edge $e$ with
$|G_{d_{0}(e)}:G_{e}|=2=|G_{d_{1}(e)}:G_{e}|$
and $G_{e}\neq 1$. Put $v=d_{0}(e)$, $w=d_{1}(e)$ and set $H=\left\langle G_{v},G_{w}\right\rangle$. Let $D=H(v\cup e\cup w)$. Since for every open subgroup $U$ of $H$ the quotient of $U\backslash D$ is connected, so is $D$. By Lemma 2.14 $H=G_{v}\amalg_{G_{e}}G_{w}$. Since $G_{e}$ is normal in $H$, by [Rib17, Proposition 4.2.2, (b)], $G_{e}$ acts trivially on $D$, contradicting the $k$-acylindricity of the action. This finishes the proof. ∎
This proposition shows that if the action is $k$-acylindrical, then we can apply the results in Subsection 2.3 without any restrictions.
Theorem 3.5.
Suppose $\mathcal{C}$ does not contain all finite groups and let $G$ be a pro-$\mathcal{C}$ group acting $k$-acylindrically on a simply connected profinite graph $\Gamma$. Then the maximal abstract subgraph $D$ of $\Gamma$ such that each $e\in D$ satisfies $G_{e}\neq 1$ has finite diameter $\leq 2k$. Moreover, for every connected component $D^{*}$ of $D$ there at most one other connected component at finite distance to $D^{*}$ in which case stabilizers of edges and vertices of $D^{*}$ are of order 22 22.
Proof.
Suppose on the contrary there exists a finite connected subgraph $\Sigma$ of $D$ whose diameter is $>2k$. Let $H=\left\langle G_{v}\mid v\in V(\Sigma)\right\rangle$ and $T=H\Sigma$. By Lemma 2.14 and [Rib17, Proposition 4.1.1], $H$ is the fundamental group of a tree of pro-$\mathcal{C}$ groups $\Pi_{1}(\mathcal{H},H\backslash T)$ whose edge and vertex groups are stabilizers of the corresponding edges and vertices of a connected transversal of $H\backslash T$ in $T$. Hence, by Proposition 3.4 there is no edge $e$ such that
$|\mathcal{H}(d_{0}(e)):\mathcal{H}(e)|=2=|\mathcal{H}(d_{1}(e)):\mathcal{H}(e)|$
with $G_{e}\neq 1$. Let $(\mathcal{H}_{red},H\backslash T_{red})$ be a reduced tree of groups obtained from $(\mathcal{H},H\backslash T)$. If it has more than one vertex, then $H$ cannot be pro-$\mathcal{C}$ by Proposition 2.12. Hence, $(\mathcal{H}_{red},H\backslash T_{red})$ has one vertex $v$ only which implies that $H$ fixes a vertex $w$ in $T$ and, in fact, $w\in\Sigma$. As the diameter of $\Sigma$ is greater than $2k$, it follows that there exists a geodesic $[u,w]$ in $\Sigma$ of length greater than $k$ which is fixed by $H_{u}\neq 1$ (see [Rib17, Corollary 4.1.6]) contradicting the $k$-acylindricity of the action. This proves the first part of the statement.
Now if there exist a connected component $D_{0}$ at finite distance from $D^{*}$ and $v\in D^{*},w\in D_{0}$ are vertices with $G_{v}\not\cong C_{2}$, then $\langle G_{v},G_{w}\rangle$ is not virtually cyclic and so by Proposition 2.12 can not be pro-$\cal C$, a contradiction. Similarly, if there exists still another connected component $D_{1}$ at finite distance from $D^{*}$ and $u\in V(D_{1})$, then $\langle G_{v},G_{u},G_{w}\rangle$ is not virtually cyclic again, contradicting Proposition 2.12 again. This finishes the proof. ∎
Corollary 3.6.
Suppose whithin the hypothesis of Theorem 3.5 $\Gamma$ has infinite diameter. Then $\Gamma$ possesses an edge $e$ with $G_{e}=1$, namely $G_{e}=1$ for any $e\not\in D$.
Corollary 3.7.
Suppose that $G\backslash D$ has finitely many connected components. Then $G$ acts on a simply connected profinite $G$-quotient graph $T$ with trivial edge stabilizers and vertex stabilizers equal to the stabilizers of some vertices $\Gamma$. In particular, $G$ has only a finite number of maximal stabilizers of vertices in $T$ up to conjugation and this number equals to the number maximal vertex stabilizers in $\Gamma$ that does not exceed the number of connected components of $G\backslash D$.
Proof.
By [CZ22, Lemma 3.2] the closure of any connected component $\overline{D^{*}}$ is a profinite subgraph of $\Gamma$ of diameter $\leq 2k$, hence $\overline{D}=\bigcup_{i}D_{i}^{*}$ has finite diameter, where $D_{i}^{*}$ runs via orbit representatives of the connected components of $D$. Moreover, as every connected component $\overline{D}_{\ast}$ of $\overline{D}$ has finite diameter, its stabilizer $\operatorname{Stab}(\overline{D}_{\ast})$ fixes some vertex $v_{\ast}\in V(\overline{D}_{\ast})$. Then collapsing all connected components of $\overline{D}$ we obtain a $G$-invariant simply connected quotient graph $T$ of $\Gamma$ (see [Rib17, Proposition 3.9.1]) on which $G$ acts with trivial edge stabilizers and only non-trivial vertex stabilizers of $T$ are stabilizers of connected component $\operatorname{Stab}(\overline{D}_{\ast})=G_{v_{\ast}}$ of $\overline{D}$. This concludes the proof. ∎
Let $G$ be the fundamental group of a finite graph of profinite groups and $S$ its standard graph. For any normal open subgroup $U$ of $G$ we set
$\tilde{U}=\left\langle U\cap H_{v}:v\in V(S)\right\rangle.$
Theorem 3.8.
Suppose $\mathcal{C}$ does not contain all finite groups and let $G$ be a pro-$\mathcal{C}$ group acting $k$-acylindrically on a simply connected profinite graph $\Gamma$ and $U$ a closed normal subgroup of $G$. Then the action of $G_{U}=G/\tilde{U}$ on $\tilde{U}\backslash\Gamma$ is $k$-acylindrical.
Proof.
By Theorem 3.5 an abstract subgraph $D$ whose edge stabilizers are non-trivial has diameter at most $k$. Consider the natural action of $G_{U}$ on $\Gamma_{U}=\tilde{U}\backslash\Gamma$. We shall prove that the same is true in $\Gamma_{U}=\tilde{U}\backslash\Gamma$; we use the subscript $U$ for images in $\tilde{U}\backslash\Gamma$.
Let $D$ be a maximal connected abstract subgraph of $\Gamma$ where each edge $e\in D$ satisfies $G_{e}\neq 1$. Note that any edge $e_{U}$ not in $D_{U}$ is an image of an edge $e\not\in D$ and since $G_{e}$ is trivial, $(G_{U})_{e_{U}}$ is trivial. Therefore, a maximal connected (abstract) subgraph of $\Gamma_{U}$ whose all edge stabilizers are non-trivial is contained in $D_{U}$ and so has diameter at most $k$, which implies that the action of $G_{U}$ on $\Gamma_{U}$ is $k$-acylindrical. ∎
4 Closed subgroups of fundamental groups of finite graphs of profinite groups with finite edge groups
In this section we will conclude an important step to prove Theorem E. The main theorem of this section give us necessary conditions to a prosoluble group be a subgroup of the fundamental group of an injective $k$-acylindrical finite graph of profinite groups with finite edge groups.
This lemma is a general version of [CZ23, Proposition 2.16]:
Proposition 4.1.
Let $G$ be the fundamental group of an injective finite graph of pro-$\mathcal{C}$ groups. Suppose that there is an edge $e$ such that $G(e)=1$ and $G(d_{i}(e))\neq 1$ for $i=0,1$. Then $G$ splits as a free profinite product.
Proof.
Let
[…]
be a slice of the graph $\Gamma$. Suppose that after remove the edge $e$ the graph $\Gamma$ is not connected:
[…]
Let $\Gamma_{1},\Gamma_{2}$ be the connected components and $G_{1},G_{2}$ its fundamental groups, respectively. They are embedded in $G$ by Lemma 2.11. It follows that $G=G_{1}\amalg G_{2}$. The other case is similar: if removing $e$ the graph remains connected
[…]
then we can choose a maximal subtree $T$ not containing $e$ and use it to calculate the fundamental group $G_{1}$ of the graph of groups restricted to $\Gamma-\{e\}$. So, $G$ will be a loop with edge correspondent to $G(e)$ and vertex correspondent to $G_{1}$ so that $G=\operatorname{HNN}(G_{1},G(e),t)=G_{1}\amalg\left\langle t\right\rangle$:
$G(e)$$G_{1}$
Now we can prove the main result of the section.
Theorem 4.2.
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups with finite edge groups, $G$ its profinite fundamental group. Suppose $\mathcal{C}$ does not contain all finite groups. If $H$ is a pro-$\mathcal{C}$ subgroup of $G$, then there exists $e\in E(\Gamma)$ and a conjugate $H_{0}$ of $H$ such that $H_{0}\cap G(e)=1$ and for any such $e$
$H_{0}\leq U_{1}\amalg U_{2}$
where $U_{1}\simeq\pi_{1}(\mathcal{U},\Delta)$ is the profinite fundamental group of an injective $k$-acylindrical finite graph of profinite groups with finite edge groups, $\Delta$ a connected component of $\Gamma\setminus\{e\}$ and $U_{2}$ is either isomorphic to $\widehat{\mathbb{Z}}$ or similarly the profinite fundamental group of an injective $k$-acylindrical finite graph of profinite groups with finite edge groups.
Proof.
Consider the action of $H$ on the standard profinite tree $S=S(G)$. By Corollary 3.6 there is an edge $e$ of $S$ for which $H_{e}$ is trivial. Then there is an open subgroup $U_{H}$ of $G$ containing $H$ such that $(U_{H})_{e}=1$. By [ZM89, Corollary 4.5] $U_{H}$ is the profinite fundamental group of a finite graph of groups $(\mathcal{U},\Delta)$ whose vertex and edge groups are $U_{H}$-stabilizers of vertices and edges of $S$. Thus one of the edge groups $U(\bar{e})$ of $(\mathcal{U},\Delta)$ is trivial. Assuming, w.l.o.g., that $(\mathcal{U},\Delta)$ is reduced we can apply Proposition 4.1 to deduce that $U_{H}$ splits as a free profinite product. As $U(\bar{e})$ is conjugate by some element $G$ in $G$ into some edge group of $(\mathcal{G},\Gamma)$ we may conjugate $H$ and $U$ by this elelment to get the required $H_{0}$. This finishes the proof. ∎
The second author [Zal24, Theorem 1.1] described the prosoluble subgroups of free profinite products. Using this we can deduce the following
Corollary 4.3.
One the following holds for $H_{0}$:
  1. (i)
    There is a prime $p$ such that $H_{0}\cap uU_{i}u^{-1}$ is pro-$p$ for every $i=1,2$ and $u\in U_{1}\amalg U_{2}$; besides if $H$ is finitely generated then $\sum_{i,u}d(H_{0}\cap uU_{i}u^{-1})\leq d(H)$.
  2. (ii)
    $H_{0}\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group.
  3. (iii)
    A conjugate of $H_{0}$ is a subgroup of $U_{i}$ for some $i=1,2$.
If $H$ is torsion-free, then item $(ii)$ of Corollary 4.3 does not occur.
It is not difficult to construct an example satisfying Theorem 4.2.
Example 4.4.
Let $G_{1},G_{2}$ be profinite Frobenius groups (or any finite Frobenius groups) with isomorphic complement $C$. Consider the group $G=G_{1}\amalg_{C}G_{2}$. Then $C$ is malnormal in $G$ and so $G$ satisfies the hypothesis of Theorem 4.1. Then any prosoluble subgroup $H$ has the structure stated in Corollary 4.3.
Proposition 4.5.
Let $G=\Pi_{1}(\mathcal{G},\Gamma)$ be the fundamental profinite group of a $k$-acylindrical finite graph of profinite groups. Let $v,w$ be vertices at distance $\geq 2k+1$. Then $\left\langle G(v),G(w)\right\rangle=G(v)\amalg G(w)$.
Proof.
Let $[v,w]$ be a shortest path between $v$ and $w$. Let $G(v,w)$ be the fundamental group of the finite graph of profinite groups restricted to $[v,w]$. By Lemma 2.11 $G(v,w)$ is a subgroup of $G$ generated by vertex groups of $[v,w]$. Let $e$ be an edge of $[v,w]$ at distance $>k$ from $w$ and $v$. Then $G(v)\cap G(e)=1=G(w)\cap G(e)$. Note that $G(v,w)$ splits over $G(e)$ as a free amalgamated profinite product $G(v,w)=G_{1}\amalg_{G(e)}G_{2}$, where $G_{1}$, $G_{2}$ are profinite groups generated by vertex groups of the connected components of $[v,w]\setminus\{e\}$ (see Lemma 2.11), so that $G(v)\leq G_{1}$ and $G(w)\leq G_{2}$. By Corollary 2.18, $\left\langle G(v),G(w)\right\rangle=G(v)\amalg G(w)$ as required. ∎
Corollary 4.6.
Let $G=\Pi_{1}(\mathcal{G},\Gamma)$ be the profinite fundamental group of a finite $k$-acylindrical graph of profinite groups. Let $G(v)$, $G(w)$ be vertex groups of vertices at distance $\geq 2k+1$ having epimorphisms $f_{v}:G(v)\to C_{p}$, $f_{w}:G(w)\to C_{q}$ to cyclic subgroups of $S_{n}$ of order $p\neq q$ respectively. Then $f_{v},f_{w}$ extend to homomorphism $f_{vw}:\left\langle G(v),G(w)\right\rangle\to S_{n}$ that in turn extends to a homomorphism $f:U\to S_{n}$ from some open subgroup $U$ containing $\left\langle G(v),G(w)\right\rangle$.
Proof.
The existence of $f_{vw}$ follows from Proposition 4.5. The existence of $f$ then follows by [RZ10, Lemma 8.3.8]. ∎
Lemma 4.7.
Let $(\mathcal{G},\Gamma)$ be a $k$-acylindrical finite graph of profinite groups, $G$ its fundamental group and $S=S(G)$ its standard profinite tree. Suppose $\mathcal{C}$ does not contain all finite groups and $H$ is a pro-$\mathcal{C}$ subgroup of $G$. Then for $v,w\in V(S)$ such that $H_{v}\neq H_{w}$ are non-trivial and not both of order 22 22, there exists an open subgroup $U$ containing $H$ such that the natural images of $v$ and $w$ in $\Delta=U\backslash S$ are at distance $\geq 2k+1$.
Proof.
By Proposition 3.4, there is no edge $e$ such that
$|G_{d_{0}(e)}:G_{e}|=2=|G_{d_{1}(e)}:G_{e}|$
[…]
Theorem 4.8.
Let $(\mathcal{G},\Gamma)$ be a $k$-acylindrical finite graph of profinite groups and $G$ its fundamental group. Let $H$ be a subgroup of $G$ such that there exist some non-trivial non-conjugate subgroups $H\cap G(v)$, $H\cap G(w)$ which are not both pro-$p$ for any prime $p$. Then $H$ is not prosoluble.
Proof.
First observe that since $H\cap G(v)$, $H\cap G(w)$ are not both pro-$p$ we can replace $H$ by its open subgroup such that $H\cap G(v)$, $H\cap G(w)$ have finite cyclic quotients $C_{p}$ and $C_{q}$ for $p\neq q$, respectively.
By Lemma 4.7 there exists an open subgroup $U$ containing $H$ such that the natural images $\bar{v}$, $\bar{w}$ of $v$ and $w$ in $\Delta=U\backslash S$ are at distance $\geq 2k+1$. Moreover, for some such $U$ the vertex groups $U(v)$ and $U(w)$ have epimorphisms to finite cyclic quotients $C_{p}$ and $C_{q}$ for $p\neq q$, respectively which are non-trivial restricting to $H\cap G(v)$ and $H\cap G(w)$ (cf. [RZ10, Lemma 8.3.8]). Choose $C_{p},C_{q}\in S_{n}$, $n>4$, such that any conjugate of them generate non-soluble subgroup (see [Jar94, Lemmas 1.1 and 1.2]) and let $f_{v}:U(v)\to C_{p}<S_{n}$, $f_{w}:U(w)\to C_{q}<S_{n}$ corresponding epimorphisms. Then by Corollary 4.6 we can choose our $U$ such that $f_{v},f_{w}$ extend to an epimorphism $f:U\to S_{n}$. Then $f$ restricted on $H$ surjects on $S_{n}$ and so $H$ is not prosoluble. ∎
Corollary 4.9.
Let $(\mathcal{G},\Gamma)$ be a $k$-acylindrical finite graph of profinite groups and $G$ its fundamental group. Let $H$ be a prosoluble subgroup of $G$. Then one of the following holds:
  1. (i)
    all non-trivial intersections $H\cap G(v)^{g}$, $g\in G,v\in V(\Gamma)$ are pro-$p$;
  2. (ii)
    $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group;
  3. (iii)
    $H\leq G(v)^{g}$ for some $g\in G,v\in V(\Gamma)$.
Proof.
[…]
Proposition 4.10.
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups, $G$ its fundamental group, $S$ its standard graph and $H$ is a prosoluble subgroup of $G$. If $H$ is finitely generated, then it has only finitely many vertex stabilizers up to conjugation. In particular, if $D$ is the maximal abstract subgraph of $\Gamma$ such that each $e\in D$ satisfies $H_{e}\neq 1$, then $D$ has finitely many connected components up to translation.
Proof.
Consider the decomposition
[…]
where $G_{U}=G/\tilde{U}$ with $\tilde{U}$ is the subgroup generated by the vertex stabilizers of $G$ and $G_{U}$ acts $k$-acylindrically on $S_{U}=\tilde{U}\backslash S$ by Theorem 3.8. Note that since $\tilde{U}$ is generated by all vertex stabilizers, then $G_{U}$ has finite stabilizers. Setting
$H_{U}=H\tilde{U}/\tilde{U}\simeq H/(\tilde{U}\cap H)$
we have the decomposition
[…]
It follows that $G_{U}$ and $H_{U}$ satisfies the hypothesis of Theorem 4.2. Besides, conjugating $H_{U}$ if necessary we may assume that $H_{U}=(H_{U})_{0}$ from Theorem 4.2 and therefore satisfies one of the itens (i), (ii) or (iii) of Corollary 4.3.
Suppose that $H_{U}$ satisfies Corollary 4.3 (i). Following the notation of Corollary 4.3, all intersections $H_{U}\cap uL_{i}u^{-1}$ are pro-$p$ for every $u$ and $L_{i}$ (we use the notation $L_{i}$ in the place of $U_{i}$ of Corollary 4.3 to avoid any confusion) and satisfies
$\sum_{i,u}d(H_{U}\cap uL_{i}u^{-1})\leq d(H_{U})\leq d(H).$
The groups $H_{U}\cap uL_{i}u^{-1}$ are finitely generated pro-$p$ groups whose induced action on $S_{U}$ is $k$-acylindrical. By [WZ18, Theorem 3.8] they decompose as a free pro-$p$ product of vertex stabilizers and a free pro-$p$ group where the number of free factors is bounded by $d(H_{U}\cap uL_{i}u^{-1})\leq d(H)$. Since the vertex stabilizers are finite, they are contained in some conjugate of $L_{i}$, that is, $(H_{U})_{v}\subset aL_{i}a^{-1}$ and the stabilizer of $v$ in $H_{U}\cap uL_{i}u^{-1}$ is
$(H_{U})_{v}\cap H_{U}\cap uL_{i}u^{-1}=(H_{U})_{v}.$
It follows that $H_{U}$ has finitely many vertex stabilizers up to conjugation for each $U$ and the number of this vertex stabilizers is bounded by the number of generators of $H$.
Now if $H_{U}$ satisfies (ii) then all vertex stabilizers are finite and conjugate (as they are Hall subgroups, cf. [RZ10, Theorem 2.3.5]).
Suppose $H_{U}$ satisfies Corollary 4.3 (iii) and $D_{U}=\{e\in S_{U}:(H_{U})_{e}\neq 1\}$. We show that $H\backslash D_{U}$ is connected. Suppose not. For any open $L$ as in Corollary 4.3 and let $D_{L}=\{e\in S_{U}\mid L_{e}\neq 1\}$. Then every connected component of $D_{U}$ is in some connected component of $D_{L}$ and $D_{U}=\bigcap_{L}D_{L}$. Hence there exists $L$ such that the number of connected components of $D_{L}$ is greater than 1 up to translation. This means the image of $D_{L}$ in $L\backslash S_{U}$ is disconnected and so $L=\coprod_{i}L_{i}\amalg F$ is a free profinite product of the fundamental groups $L_{i}=\Pi_{1}(\mathcal{L},\Delta_{i})$, where $\Delta_{i}$ are connected components of $L\backslash D_{L}$. Then one can write $L=L_{1}\amalg L_{2}$ with $L_{i}\leq L_{1}$, $L_{j}\leq L_{2}$ for some $i\neq j$. This means that there are $v\neq w\in S_{U}$ such that $H_{v}\leq U_{1}$ and $H_{w}\leq U_{2}$, contradicting the assumption Corollary 4.3 (iii). Thus $D_{U}$ has just one connected component up to translation in this case.
By Corollary 3.7 $H_{U}$ has a finite number of maximal vertex stabilizers up to conjugation not exceeding $d(H)$ and therefore the number of maximal vertex stabilizers of $H$ acting on $S$ does no exceed $d(H)$. Then the number of the connected components of $D$ up to translation does not exceed $d(H)$ as well. ∎
5 The finitely generated case
In this section we shall study finitely generated prosoluble subgroups of profinite groups acting $k$-acylindrically on profinite trees.
Theorem F.
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups, $G$ its fundamental group. If $H$ is a finitely generated prosoluble subgroup of $G$, then $H$ embeds into a free profinite product $\coprod_{v\in\Gamma}H_{v}$ of finitely many finitely generated profinite groups $H_{v}=H\cap\mathcal{G}(v)^{g}$ and $g\in G$. Moreover, if $H\neq H_{v}$ for some $v$, then either $H_{v}$ are finitely generated pro-$p$ or $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group (i.e., $C$ is cyclic, $|C|$ divides $p-1$ for any $p\in\pi$ and $[c,z]\neq 1$ for all $1\neq c\in C,0\neq z\in\widehat{\mathbb{Z}}_{\pi}$).
Proof.
Note that the abstract subgraph $D=\{m\mid H_{m}\neq 1\}$ has finite diameter by Theorem 3.5, and by Proposition 4.10 $D$ has finitely many connected components up to translation. Therefore $H\backslash D$ has finitely many connected components and so by Corollary 3.7 there exists a simply connected $H$-quotient graph $\overline{S}$ on which $H$ acts with trivial edge stabilizers.
By item (i) of Theorem 2.23, $(H,V(\overline{S}))$ is a $\mathcal{S}$-projective pile where $\mathcal{S}$ is the class of all finite soluble groups. By [Zal24, Remark 4.9] there exists a second countable space $T$ with the same non-trivial point stabilizers such that $(H,T)$ is a projective pile. Since $T$ is second countable, the quotient map $T\to H\backslash T$ has a continuous section (see [RZ10, Lemma 5.6.7] or [Rib17, 1.3.10]). By Theorem 2.23 $(H,T)$ is a $\mathcal{S}$-projective pile such that the quotient map $T\to H\backslash T$ has a continuous section. Therefore by Proposition 2.24 $H$ embeds into a free prosoluble product
[…]
where $F$ is a free prosoluble group. Moreover, every $H_{t}$ is of the form $H\cap G(v)^{g}$. By Corollary 4.9 $H_{t}$ are either all pro-$p$ or $H$ is Frobenius and all $H_{t}$ are conjugate as they are Hall subgroups of $H$. Then by [Zal24, Theorem 1.4] $H$ embeds into a desired free profinite product. ∎
Corollary 5.1.
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups, $G$ its fundamental group and $S$ its standard profinite graph. If $H$ is a finitely generated prosoluble subgroup of $G$, then it has only finitely many vertex stabilizers up to conjugation. Moreover, $\sum_{v}d(H_{v})\leq d(H)$, where $v$ runs through representatives of $H$-orbits in $S$.
Proof.
Follows from Theorem F and [Zal24, Corollary 1.6]. ∎
6 Hyperbolic virtually special groups
In this section we apply the previous results to hyperbolic and relatively hyperbolic virtually compact special groups.
We start with some definitions.
Definition 6.1.
A group is virtually special if there is a finite index subgroup isomorphic to the fundamental group of a compact special cube complex. A group is virtually compact special if there is a finite index subgroup isomorphic to the fundamental group of a compact special cube complex [HW08, Definition 3.2].
Definition 6.2.
A family $\{H_{1},...,H_{n}\}$ of subgroups of $G$ is malnormal if, for any $g\in G$, $H_{i}\cap H_{j}^{g}\neq 1$ implies $i=j$ and $g\in H_{i}$.
Definition 6.3.
A hierarchy of groups of length 00 0 is a single vertex labeled by a group. A hierarchy of groups of length $n$ is a graph of groups $(\mathcal{G}_{n},\Gamma_{n})$ together with hierarchies of length $n-1$ on each vertex of $\Gamma_{n}$ (cf. [WZ17, Definition 1.3]).
In [WZ17, Theorem 1.4] it is shown that a hyperbolic group $G$ is virtually compact special if, and only if, $G$ has a subgroup $\Gamma_{0}$ of finite index with a malnormal hierarchy. It is also shown that, in this case, $\Gamma_{0}$ has a 11 11-acylindrical action on a profinite tree (see [WZ17, Lemma 7.3]). In the relatively hyperbolic virtually compact special case, $\Gamma_{0}$, in addition, has a malnormal hierarchy terminating in parabolic subgroups (see [Zal23a, Theorem 3.2]).
Theorem B.
Let $G$ be a torsion-free, hyperbolic, virtually compact special group. Any finitely generated prosoluble subgroup $H$ of the profinite completion of $G$ is projective.
Proof.
Let $G$ be a finitely generated hyperbolic virtually compact special group and $H$ a finitely generated prosoluble subgroup of $\widehat{G}$. Let $H_{0}=\widehat{\Gamma}_{0}\cap H$ where $\Gamma_{0}$ is the subgroup defined above. We use induction on the length of the malnormal hierarchy described above. Since $\widehat{\Gamma_{0}}$ is a fundamental group of a graph of profinite groups whose action on the standard graph is 11 11-acylindrical (see [WZ17, Lemma 7.3]), by Theorem F, we have an embedding
$H_{0}\hookrightarrow\coprod H_{0}\cap G(v)^{g}.$
Then by Corollary 4.3 either $H_{0}\leq H_{v}^{g}$ or all $H_{0}\cap G(v)^{g}$ are pro-$p$. In the first case the result follows immediately from the induction hypothesis. In the second case we observe first that, by Theorem F, $H_{0}\cap G(v)^{g}$ are finitely generated which allows to apply (cf. [WZ17, Theorem F]) to deduce that $H_{0}\cap G(v)^{g}$ are free pro-$p$. It follows that $H_{0}$ is projective and so $H$ is virtually projective. Since it is torsion-free, it is projective (see [Zal04, Proposition 2.5]). ∎
Theorem B yields a description of prosoluble subgroups of closed hyperbolic 33 33-manifolds.
Theorem A.
Let $\pi_{1}(M)$ be the fundamental group of a compact hyperbolic 33 33-manifold $M$ and $H$ a finitely generated prosoluble subgroup of the profinite completion $\pi_{1}(M)$. Then $H$ is projective.
Proof.
By [AFW15, Theorem 5.4] $\pi_{1}(M)$ is virtually compact special. Hence the result follows from Theorem B. ∎
Let $G$ be a residually finite group hyperbolic relative to a malnormal family of parabolic subgroups $\mathcal{P}=\{P_{1},...,P_{n}\}$. A closed subgroup of the profinite completion of some $P_{i}$ will be called parabolic subgroup of $\widehat{G}$.
Proposition 6.4.
Let $G$ be a relatively hyperbolic virtually compact special group and $H$ a finitely generated prosoluble subgroup of the profinite completion of $G$. Then there is an open subgroup $H_{0}$ of $H$ such that one of the following statements holds:
  1. (i)
    $H_{0}$ is isomorphic to a subgroup of a free prosoluble product of pro-$p$ subgroups of parabolic subgroups,
  2. (ii)
    $H_{0}$ is a subgroup of a parabolic subgroup.
Proof.
Let $G$ be a finitely generated relatively hyperbolic virtually compact special group and $H$ a finitely generated prosoluble subgroup of $\widehat{G}$. In [Zal23a, Proof of Theorem 1.3] we see that there is a subgroup $\Gamma_{0}$ such that $\widehat{\Gamma_{0}}$ acts 11 11-acylindrically on a profinite tree. Let $H_{0}=\widehat{\Gamma}_{0}\cap H$. As in the proof of Theorem B, by Theorem F, we have an embedding
$H_{0}\hookrightarrow\coprod H_{0}\cap G(v)^{g}.$
We assume w.l.o.g that $H_{0}$ is torsion-free. Then by Corollary 4.3 either $H_{0}\leq H_{v}^{g}$ or all $H_{0}\cap G(v)^{g}$ are pro-$p$. In the first case the result follows immediately from the induction hypothesis.
In the second case we observe first that by Corollary 4.10 $H_{0}\cap G(v)^{g}$ are finitely generated which allows to apply (cf. [Zal23a, Theorem 1.4]) to deduce that $H_{0}\cap G(v)^{g}$ are free pro-$p$ products of parabolic pro-$p$ groups and a free pro-$p$ group. ∎
Definition 6.5.
We say that a relatively hyperbolic group is toral if it is torsion-free hyperbolic relative to a set $\{A_{1},...,A_{n}\}$ of parabolic subgroups where each $A_{i}$ is a finitely generated abelian group.
Corollary 6.6.
Let $G$ be a torsion-free relatively hyperbolic virtually compact special group and $H$ a finitely generated prosoluble subgroup of the profinite completion $\widehat{G}$ of $G$. Then $H$ is virtually abelian or virtually a subgroup of a free prosoluble product of abelian pro-$p$ groups.
7 The fundamental group of standard compact arithmetic manifolds
In this section we give a direct application of Proposition 6.4.
Let $\operatorname{SO}(n,1)$ be the special orthogonal group and $R_{k}$ the ring of integers of an algebraic number field $k$. Recall that a hyperbolic $n$-manifold is a manifold $M^{n}=\mathbb{H}^{n}/\Gamma$ such that $\Gamma$ is a torsion-free subgroup of $O(n,1)$.
Definition 7.1.
A standard arithmetic subgroup $\Gamma$ of $\operatorname{SO}(n,1)$ is a group commensurable with the subgroup $\operatorname{SO}(q,R_{k})$ of $R_{k}$-points of
$\operatorname{SO}(q)=\{X\in\operatorname{SL}(n+1,\mathbb{C}):X^{t}qX=q\}.$
An arithmetic hyperbolic $n$-manifold $M^{n}=\mathbb{H}^{n}/\Gamma$ such that $\Gamma$ is standard arithmetic is a standard arithmetic manifold.
It is shown in [BHW11] that standard uniform arithmetic subgroups of $\operatorname{SO}(n,1)$ are virtually compact special. Thus, Theorem B can be applied to standard arithmetic hyperbolic manifolds. Also, it is well-known that the fundamental group of these manifolds is hyperbolic relative to virtually abelian subgroups. Therefore, Proposition 6.4 yields
Theorem D.
Let $M$ be a standard uniform hyperbolic arithmetic manifold and $H$ a finitely generated prosoluble subgroup of the profinite completion of $\pi_{1}(M)$. Then $H$ is projective.
8 The fundamental group of compact 33 33-manifolds
To prove Theorem E we will use the pro-$p$ version of it, i.e., the classification of pro-$p$ subgroups of the profinite completion of a 3-manifold group established by Henry Wilton and the second author in [WZ18, Theorem 1.3].
Theorem 8.1.
[WZ18, Theorem 1.3] A finitely generated pro-$p$ subgroup of the profinite completion of the fundamental group of a compact, orientable 33 33-manifold $M$ is a free pro-$p$ product of the pro-$p$ groups from the following list of isomorphism types:
  1. (i)
    For $p>3$: $C_{p}$; $\mathbb{Z}_{p}$; $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$; the pro-$p$ completion of $(\mathbb{Z}\times\mathbb{Z})\rtimes\mathbb{Z}$ and the pro-$p$ completion of a residually-$p$ fundamental group of a non-compact Seifert fibred manifold with hyperbolic base of orbifold;
  2. (ii)
    For $p=3$: in addition to the list of (i) we have a torsion-free extension of $\mathbb{Z}_{3}\times\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ by $C_{3}$;
  3. (iii)
    For $p=2$: in addition to the list of (i) we have $C_{2^{m}}$; $D_{2^{k}}$; $Q_{2^{n}}$; $\mathbb{Z}_{2}\rtimes C_{2}$; the torsion-free extensions of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ by one of $C_{2}$, $C_{4}$, $C_{8}$, $D_{2}$, $D_{4}$, $D_{8}$, $Q_{16}$; the pro-22 22 extension of the Klein-bottle group $\mathbb{Z}\rtimes\mathbb{Z}$; the pro-22 22 completion of all torsion-free extensions of a soluble group $(\mathbb{Z}\rtimes\mathbb{Z})\rtimes\mathbb{Z}$ with a group of order at most 22 22.
Now we are ready to prove the theorem that classifies prosoluble subgroups of the profinite completion of 3-manifold groups.
Theorem E.
Let $M$ be a compact orientable 33 33-manifold. If $H$ is a finitely generated prosoluble subgroup of $\widehat{\pi_{1}(M)}$, then one of the following statements holds:
  1. (i)
    $H$ is isomorphic to a subgroup of a free prosoluble product of the pro-$p$ groups from the following list of isomorphism types:
    1. (1)
      For $p>3$: $C_{p}$; $\mathbb{Z}_{p}$; $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$; the pro-$p$ completion of $(\mathbb{Z}\times\mathbb{Z})\rtimes\mathbb{Z}$ and the pro-$p$ completion of a residually-$p$ fundamental group of a non-compact Seifert fibred manifold with hyperbolic base of orbifold;
    2. (2)
      For $p=3$: in addition to the list of (1) we have a torsion-free extension of $\mathbb{Z}_{3}\times\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ by $C_{3}$;
    3. (3)
[…]
  2. (ii)
    $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group where $C$ is a finite cyclic group and $\pi$ is set of primes;
  3. (iii)
    $H$ is a subgroup of a central extension of either $D_{2n}$ or tetrahedral group $T$ or octahedral group $O$ or a icosahedral group $I$ by a cyclic group of even order;
  4. (iv)
    $H$ is a subgroup of the profinite completion of a 33 33-dimensional Bieberbach group $B$ (i.e, $B$ is torsion-free virtually $\mathbb{Z}^{3}$);
  5. (v)
    $H$ is a subgroup of the profinite completion of a group extension of $\mathbb{Z}^{2}\rtimes_{A}\mathbb{Z}$ by $C$ where $C$ is trivial or $C_{2}$ and $A$ is an Anosov matrix.
  6. (vi)
    $H$ is an extension of a torsion-free procyclic group ^$\widehat{\mathbb{Z}}_{\sigma}$ by a subgroup $H_{0}$ of a finite free prosoluble product of finite cyclic $p$-groups with $H_{0}$ acting either trivially on ^$\widehat{\mathbb{Z}}_{\sigma}$ or by inversion.
8.1 Reduction to the irreducible case
Let $M$ be a compact, orientable, non-geometric 33 33-manifold. Since every compact manifold is a retract of a closed manifold, it follows easily that we may reduce to the case in which $M$ is closed.
If $M$ is reducible, by the Prime Decomposition Theorem, we can write
$\pi_{1}(M)=\pi_{1}(M_{1})\ast\cdots\ast\pi_{1}(M_{n})$
where each $M_{i}$ is prime 33 33-manifold. A prime 33 33-manifold is either irreducible or $S^{2}\times S^{1}$ or the non-orientable $S^{2}$ bundle over $S^{1}$. In the last two cases, the fundamental group is isomorphic to $\mathbb{Z}$ (see [Hem04] and [GSW21, Page 5]).
Thus applying [Zal24, Theorem 1.4] we obtain
Theorem 8.2.
Let $H$ be a finitely generated prosoluble subgroup of the profinite completion $G=\widehat{\pi_{1}(M)}$ of a reducible 33 33-manifold having non-trivial free product decomposition
$\pi_{1}(M)=\pi_{1}(M_{1})\ast\cdots\ast\pi_{1}(M_{n})\ast F$
where each $M_{i}$ is irreducible 33 33-manifold and $F$ is a free group of finite rank. Then one of the following holds:
  1. (i)
    all intersection $H\cap\widehat{\pi_{1}(M_{i})}^{g}$, $g\in G$ are pro-$p$ for some fixed prime $p$ and $\sum_{v}d(H\cap\widehat{\pi_{1}(M_{i})}^{g})\leq d(H)$;
  2. (ii)
    $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group (i.e $C$ is cyclic, $|C|$ divides $p-1$ for any $p\in\pi$ and $[c,z]\neq 1$ for all $1\neq c\in C,0\neq z\in\widehat{\mathbb{Z}}_{\pi}$);
  3. (iii)
    $H\leq\widehat{\pi_{1}(M_{i})}^{g}$ for some $g\in G$.
In case $(i)$ we apply Theorem 8.1 to $H\cap\widehat{\pi_{1}(M_{i})}^{g}$ to deduce the structure of each of these intersections and obtain $(i)$ of Theorem E. Case $(ii)$ of the theorem above corresponds to $(ii)$ of Theorem E.
Thus we only need to consider $(iii)$ and in this case we may assume, w.l.o.g, that $H\leq\widehat{\pi_{1}(M_{i})}$, i.e., we may assume that $M$ is irreducible.
8.2 Reduction to the geometric cases
Now we can use the following
Proposition 8.3 ([WZ18], Proposition 4.1).
Let $M$ be a closed irreducible 33 33-manifold. Then the action of $\widehat{\pi_{1}(M)}$ on the standard profinite tree $S$ associated to the JSJ-decomposition of $M$ is 44 44-acylindrical.
Then we can apply Theorem F to deduce
Theorem 8.4.
Let $M$ be a closed 33 33-manifold. If $H$ is a finitely generated prosoluble subgroup of $\widehat{\pi_{1}(M)}$, then $H$ embeds into a free profinite product $\coprod_{v\in V}H_{v}$ of $H$-stabilizers of vertices in $S$ and one of the following statements holds:
  1. (i)
    $H$ is isomorphic to a subgroup of a free prosoluble product of finitely many finitely generated pro-$p$ groups $H\cap uH_{v}u^{-1}$ for every $v\in V$ and $u\in\coprod_{v\in V}H_{v}$;
  2. (ii)
    $H$ is a profinite Frobenius group ^$\widehat{\mathbb{Z}}_{\pi}\rtimes C$;
  3. (iii)
    $H$ is a subgroup of $H_{v}$ for some $v\in V$.
If $H$ satisfies $(i)$, then we can apply Theorem 8.1 again and deduce $(i)$ or $(ii)$ of Theorem E. The case $(ii)$ corresponds to Theorem E $(ii)$. Therefore we may assume that $H$ satisfies item $(iii)$ and since $H$ in this case is conjugate into a vertex of $\widehat{\pi_{1}M}$ we may assume that $H$ is contained in the profinite completion of a geometric manifold.
Thus, to complete the proof of Theorem E we need to study the profinite completion of $\pi_{1}(M)$ when $M$ is geometric.
8.3 Geometric cases
The Geometrization conjecture, proved by G. Perelman on his approach to prove the Poincaré conjecture (see [Per02], [Per03b], [Per03a]), says that a 33 33-manifold is geometrically modeled by one of the eight Thurston’s geometries
  1. (i)
    $S^{3}$,
  2. (ii)
    $S^{2}\times\mathbb{R}$,
  3. (iii)
    $\mathbb{R}^{3}$,
  4. (iv)
    $Nil$,
  5. (v)
    $Sol$
  6. (vi)
    $\mathbb{H}^{3}$,
  7. (vii)
    $\mathbb{H}^{2}\times\mathbb{R}$,
  8. (viii)
    $\widetilde{\operatorname{SL}_{2}(\mathbb{R})}$.
The fundamental group of the first five are, respectively, of the following types:
  1. (i)
    a finite cyclic group or a central extension of $D_{2n}$, $T$, $O$ or $I$ by a cyclic group of even order. We are denoting by $T$ the tetrahedral group of 1212 1212 rotational symmetries of a tetrahedron, by $O$ the octahedral group of 2424 2424 rotational symmetries of a cube or of an octahedron, by $I$ the icosahedral group of 6060 6060 rotational symmetries of a dodecahedron or an icosahedron;
  2. (ii)
    isomorphic to $\mathbb{Z}$ or the infinite dihedral group;
  3. (iii)
    virtually abelian;
  4. (iv)
    virtually nilpotent;
  5. (v)
    virtually soluble.
The fundamental groups of these first five geometries are soluble. If we consider $H$ as a prosoluble subgroup of the profinite completion of these groups, then $H$ can be any of its subgroups. We summarize the existing classification as:
  1. (i)
    $H$ is a finite cyclic group or $H$ is a subgroup of a central extension of a dihedral, tetrahedral, octahedral or icosahedral group by a cyclic group of even order;
  2. (ii)
    $H$ is a subgroup of the profinite completion of the infinite dihedral group;
  3. (iii)
    $H$ is a subgroup of the profinite completion of a 33 33-dimensional Bieberbach group (i.e, virtually $\mathbb{Z}^{3}$);
  4. (iv)
    $H$ is a subgroup of the profinite completion of a group extension of $\mathbb{Z}^{2}\rtimes_{A}\mathbb{Z}$ by $C$ where $C$ is trivial or isomorphic to $C_{2}$ and $A$ is an Anosov matrix;
The reader can check [AFW15], [Sco83, Section 4], [Ner21, Section 1.5, Chapter 2, Chapter 3] and [GM23] for details on the fundamental groups of the first five items.
The item $(vi)$ (hyperbolic case) is the subject of Theorem A. Since a finitely generated projective profinite group is a subgroup of a non dihedral free profinite product (cf. Proposition 2.12) it is a particular case of Theorem E, item $(i)$. So, it remains to deal with cases $(vii)$ and $(viii)$.
8.4 Seifert manifolds
In order to conclude the analysis of the possible irreducible compact 33 33-manifolds we now deal with the Seifert manifolds. We have the following exact sequence:
$1\to\mathbb{Z}\to\pi_{1}(M)\to\pi_{1}(O)\to 1$
where $\pi_{1}(O)$ is a Fuchsian group acting on $\mathbb{Z}$. Thus we describe first prosoluble subgroups of a profinite completion of a Fuchsian group.
Theorem 8.5.
Let $H$ be a prosoluble subgroup of a profinite completion of a Fuchsian group $\widehat{\pi_{1}(O)}$. Then one of the following holds:
  1. (i)
    $H$ is a subgroup of a finite free prosoluble product of finite cyclic $p$-groups;
  2. (ii)
    $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group;
  3. (iii)
    $H$ is a finite cyclic group.
Proof.
Passing to an open subgroup containing $H$ if necessary (see [HKS71]), we can replace $O$ by a non-trivial finite-sheeted cover of degree $n>2$ whose profinite completion contains $H$, hence, we can assume that $\pi_{1}(O)$ is not a triangle group. In this case we have
$\widehat{\pi_{1}(O)}=\widehat{K}\amalg_{\widehat{C}}\widehat{F}$
where $K$ is a free product of cyclic groups, $F$ a free profinite group of finite rank and $C$ an infinite cyclic group (see [LMR96]).
Since $\widehat{C}$ is malnormal in $\widehat{\pi_{1}(O)}$, by [WZ17, Lemma 7.3], $\widehat{\pi_{1}(O)}$ is a 11 11-acylindrical finite graph of profinite groups. Applying Theorem F we get one of the following cases:
  • (a)
    $H$ is subgroup of the free prosoluble product of $H\cap\widehat{F}^{g}$ or $H\cap\widehat{K}^{g}$ and these intersection are pro-$p$.
  • (b)
    $H$ is a profinite Frobenius group.
  • (c)
    $H$ is contained in either $\widehat{F}$ or $\widehat{K}$ up to conjugation.
Suppose that (a) holds. Note that $H\cap\widehat{F}^{g}$ is free pro-$p$ and $H\cap\widehat{K}^{g}$ is free pro-$p$ product of cyclic pro-$p$ groups by the pro-$p$ version of the Kurosh Subgroup Theorem (cf. [Rib17, Theorem 9.6.2]). Then we have $(i)$. If $H$ is Frobenius, we have item $(ii)$.
Suppose that (c) holds. If $H$ is the conjugate of $\widehat{F}$ then it is projective and so $(i)$ holds. Otherwise we may assume that $H\leq\widehat{K}$. Then applying Theorem F again we deduce the statement of the theorem. ∎
Theorem 8.6.
Let $G$ be the fundamental group of a Seifert manifold with either geometry $\mathbb{H}^{2}\times\mathbb{R}$ or $\widetilde{\operatorname{SL}_{2}(\mathbb{R})}$ and $H$ a finitely generated prosoluble subgroup of $\widehat{G}$. Then $H$ is an extension of a torsion-free procyclic group ^$\widehat{\mathbb{Z}}_{\sigma}$ by a subgroup of a finite free prosoluble product of finite cyclic $p$-groups.
Proof.
We have the following exact sequence:
$1\to\mathbb{Z}\to\pi_{1}(M)\to\pi_{1}(O)\to 1$
where $\pi_{1}(O)$ is a Fuchsian group acting on $\mathbb{Z}$ either trivially or by inversion. Note that we have an action of $\widehat{\pi_{1}(O)}$ therefore of $H_{0}=H\widehat{\mathbb{Z}}/\widehat{\mathbb{Z}}$ on $\widehat{\mathbb{Z}}$ induced by the action of $\pi_{1}(O)$ on $\mathbb{Z}$ which is trivial or by inversion.
Thus we have an exact sequence
$1\to\widehat{\mathbb{Z}}_{\sigma}\to H\to H_{0}\to 1,$
where ^$\widehat{\mathbb{Z}}_{\sigma}=H\cap\widehat{\mathbb{Z}}$ and $H_{0}$ acts on ^$\widehat{\mathbb{Z}}_{\sigma}$ either trivially or by inversion.
By Theorem 8.5 we have the following 3 possibilities for $H_{0}$.
  1. (a)
    $H_{0}$ is a subgroup of a finite free prosoluble product of finite cyclic $p$-groups;
  2. (b)
    $H_{0}\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group;
  3. (c)
    $H_{0}$ is a finite cyclic group.
and we analyze each of them in turn.
  1. (a)
    In this case $H$ is an extension of ^$\widehat{\mathbb{Z}}_{\sigma}$ by $H_{0}$ with the either trivial action or by inversion.
  2. (b)
    Suppose that
    $1\to\widehat{\mathbb{Z}}\to H\to\widehat{\mathbb{Z}}_{\pi}\rtimes C_{n}\to 1$
    is an exact sequence with ^$\widehat{\mathbb{Z}}_{\pi}\rtimes C_{n}$ a profinite Frobenius group with $C_{n}$ a finite cyclic group of order $n$. As the lift of every element of finite order of $\pi_{1}(O)$ to $\pi_{1}(M)$ must centralize $\mathbb{Z}$, the lift of every element of finite order of $H_{0}$ must centralize ^$\widehat{\mathbb{Z}}_{\sigma}$. But $H_{0}$ is generated by torsion elements, so the extension is central in this case.
    Let $\rho$ be the set of all primes dividing $n$. Then the preimage of $C_{n}$ in $H$ is ^$\widehat{\mathbb{Z}}_{\rho^{\prime}}\times\widehat{\mathbb{Z}}_{\rho}$ where $\rho^{\prime}$ is the complement of $\rho$ in $\sigma$ (it is isomorphic but not equal to ^$\widehat{\mathbb{Z}}_{\sigma}$. Thus we can write $H$ as (^$(\widehat{\mathbb{Z}}_{\rho^{\prime}}\times\widehat{\mathbb{Z}}_{\pi})\rtimes\widehat{\mathbb{Z}}_{\rho}\cong\widehat{\mathbb{Z}}_{\rho^{\prime}}\rtimes(\widehat{\mathbb{Z}}_{\pi}\rtimes\widehat{\mathbb{Z}}_{\rho})$. As $\pi\cap\rho=\emptyset$, ^$\widehat{\mathbb{Z}}_{\pi}\rtimes\widehat{\mathbb{Z}}_{\rho}$ is projective and hence a subgroup of a free profinite product of cyclic $p$-groups.
  3. (c)
    Suppose that
    $1\to\widehat{\mathbb{Z}}_{\sigma}\to H\to C_{n}\to 1$
    is an exact sequence with $C_{n}$ finite cyclic of order $n$. As the lift of every element of finite order of $\pi_{1}(O)$ to $\pi_{1}(M)$ must centralize $\mathbb{Z}$, the lift of every element of finite order of $H_{0}$ must centralize ^$\widehat{\mathbb{Z}}_{\sigma}$. Hence the extension is central and so $H$ is procyclic torsion-free.